What is the probability that 6 rolls of a fair die will show exactly three fives?
The probability of getting three fives in the first three rolls and non-fives
in the
next three rolls is; P(5,5,5,N5,N5,N5) =
1/6 x 1/6 x 1/6 x 5/6 x 5/6 x 5/6
=
0.002679...
The number of different order in which the fives can come out is given by;
6C3 =
6!/[3!∙(6-3)!] =
20
So the probability that in 6 rolls of a fair die exactly three fives (in any order) will
come out is; P(three
fives any order) =
(20)∙(1/6)3∙(5/6)3 =
0.05358... ~ 5.4%
1/16
When a coin is tossed 7 times in a row what is the probability that heads comes up?
The chance is 3.5/7.
What is the probability that the total will be 8 when you roll two dices?
the answer is 5/36 coz (2,6) (3,5) (4,4) (5,3) (6,2) which are the five possible outcomes out of the 36 total no. of outcomes. Enjoy!! Kush
If two dice are rolled one time what is the probability of getting the number 8?
Sum of two dice eqauls 8:
2,6
3,5
4,4
5,3
6,2
= 5 in 36
How do you do probability in math?
The Probability of Success = Number of successful outcomes/Number of outcomes.
E.g. Find the probability of choosing a red five or a black odd numbered card in a standard deck of 52 playing cards.
There are 2 red fives and 10 odd numbered cards, and a total of 52 cards, so:
=2/52+10/52
=12/52
3/13
Probability of Red 5 or a Black Odd = 3/13
* * * * *
The above is true only for discrete distributions, not for continuous variables. For a continuous variable, with probability distribution function p(x), the probability that x lies between two values, a and b (ie a
What does The first law of probability state?
States that the results of one chance event have no effect on the results of subsequent chance events.
What is the probability of getting heads only once if a fair coin is tossed 3 times?
If heads comes up on the first throw (p=0.5) then tails must come up on the next two throws (each has a p= 0.5 of occurring) so 0.5 x 0.5 x 0.5 = 0.125 for this event (HTT) The second event is THT and the third event TTH each has a 0.125 probability, so 3 x 0.125 = 0.375. You can also solve this problem with the binomial distribution p = 0.5 q= 0.5, p(x=1, n= 3) = 3!/(2! 1!)(0.5)^1(0.5)^2 = 0.375 If heads comes up on the first throw (p=0.5) then tails must come up on the next two throws (each has a p= 0.5 of occurring) so 0.5 x 0.5 x 0.5 = 0.125 for this event (HTT) The second event is THT and the third event TTH each has a 0.125 probability, so 3 x 0.125 = 0.375. You can also solve this problem with the binomial distribution p = 0.5 q= 0.5, p(x=1, n= 3) = 3!/(2! 1!)(0.5)^1(0.5)^2 = 0.375
What is the probability of drawing a king of hearts from a regular deck of 52 cards?
(question not clear) , as far to my understanding it is 4 over 52 simplified ,
so answer is = 1 over 13. (probability of a king drawn from a pack of cards)
How do you calculate expected value for a probability situation?
simulate 100 events where 12 students have been accepted and each student is 66% likely to accept. andministrators want to try to ubderstand how many students will attend the school
If there are 5 questions, each with 2 possible answers, there will be 25 different quiz outcomes possible. We know that 25 = 32 different possible outcomes on a competed quiz where one answer is selected for each question.
If DNA test says 0 probability can it be wrong?
Only if a mistake happened in the lab. This could be due to a mislabeled tube a sample placed in the wrong tube or any number of other misplaced samples or data. Most labs have systems in place to minimise these errors but none are perfect. Humans carry out these tests and humans make mistakes.
If a coin is tossed for 30 times what is the probability of having no heads?
No heads means that every toss lands tails. (0.5)30= 9.3 x 10-10 Note that 109 = 1 trillion, so the probability can be stated this event is likely to occur about 9 times in 10 trillion tosses.
What is the probability of rolling 2 dice whose sum is a prime?
All possible outcomes: 2,3,4,5,6,7,8,9,10,11,12
All possible outcomes that are prime: 2,3,5,7,11
All possible rolls that are one of these prime numbers:
--1-2-3-4-5-6
1.*.*....*....*
2.*....*....*...
3....*....*......
4.*....*.........
5....*..........*
6.*..........*...
So we count 15 possible ways to make a prime number.
P = (# of correct outcomes) / (# of total possible outcomes)
P = 15 / (6*6) = 0.416667 or 5/12
3 out of 6
Probability of precipitation means what?
Forecasts issued by the National Weather Service routinely include a "PoP" (probability of precipitation) statement, which is often expressed as the "chance of rain" or "chance of precipitation".
EXAMPLEZONE FORECASTS FOR NORTH AND CENTRAL GEORGIA
NATIONAL WEATHER SERVICE PEACHTREE CITY GA
119 PM EDT THU MAY 8 2008
GAZ021-022-032034-044046-055-057-090815-
CHEROKEE-CLAYTON-COBB-DEKALB-FORSYTH-GWINNETT-HENRY-NORTH FULTON-
ROCKDALE-SOUTH FULTON-
INCLUDING THE CITIES OF...ATLANTA...CONYERS...DECATUR...
EAST POINT...LAWRENCEVILLE...MARIETTA
119 PM EDT THU MAY x 2008
.THIS AFTERNOON...MOSTLY CLOUDY WITH A 40 PERCENT CHANCE OF
SHOWERS AND THUNDERSTORMS. WINDY. HIGHS IN THE LOWER 80S. NEAR
STEADY TEMPERATURE IN THE LOWER 80S. SOUTH WINDS 15 TO 25 MPH.
.TONIGHT...MOSTLY CLOUDY WITH A CHANCE OF SHOWERS AND
THUNDERSTORMS IN THE EVENING...THEN A SLIGHT CHANCE OF SHOWERS
AND THUNDERSTORMS AFTER MIDNIGHT. LOWS IN THE MID 60S. SOUTHWEST
WINDS 5 TO 15 MPH. CHANCE OF RAIN 40 PERCENT.
What does this "40 percent" mean? ...will it rain 40 percent of of the time? ...will it rain over 40 percent of the area?
The "Probability of Precipitation" (PoP) describes the chance of precipitation occurring at any point you select in the area.
How do forecasters arrive at this value?
Mathematically, PoP is defined as follows: PoP = C x A where "C" = the confidence that precipitation will occur somewhere in the forecast area, and where "A" = the percent of the area that will receive measureable precipitation, if it occurs at all.
So... in the case of the forecast above, if the forecaster knows precipitation is sure to occur ( confidence is 100% ), he/she is expressing how much of the area will receive measurable rain. ( PoP = "C" x "A" or "1" times ".4" which equals .4 or 40%.)
But, most of the time, the forecaster is expressing a combination of degree of confidence and areal coverage. If the forecaster is only 50% sure that precipitation will occur, and expects that, if it does occur, it will produce measurable rain over about 80 percent of the area, the PoP (chance of rain) is 40%. ( PoP = .5 x .8 which equals .4 or 40%. )
In either event, there is a correct way to interpret the forecast: According to the NWS, "CHANCE OF RAIN 40 PERCENT" means there is a 40 percent chance that rain will occur at any given point in the area.
NOTE! However, this is not entirely accurate language. It is better expressed as: "CHANCE OF RAIN 40 PERCENT" means the on average for all of the points in the area during the specified time period (usually 12 hour periods), chance that rain will occur (on average) is 40%.
Explanation:Suppose the the forecast were for Maui, HI. One "given" point is your house near the top of Mt. Haleakala, where it rains almost constantly. 40% percent is obviously not accurate for that given point. So assume that Mt. Haleakala is 10% of the area of Maui and that the average chance of rain today there is 80%. And assume that the average chance of rain for the other 90% of the island is 35%. So for the entire island, the average chance of rain is (.9 * .35) + (.1 * .8) = .4 = 40%.Clearly, the smaller the area, the more meaningful and accurate "chance of rain" is.
A spinner has the numbers 1 thru 9.
What is the probability of P(less than 6)? Write the answer as a decimal.